Please help me solve this. If you have a recipe that calls for 1 1/4 cups of water to 3 cups of flour, How many cups of water would you need for 2 cups of flour.
I have tried using 5/4 for the water and multiplying and can't seen to get the answer.

Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

1 1/4 cups of water to 3 cups of flour
5/4 cups of water to 3 cups of flour
Divide both sides by 3
5/12 of a cup of water to 1 cup of flour
Multiply both sides by 2
5/6 of a cup of water to 2 cups of flour.
Hope this helps.
:-)

Answer by stanbon(75887)

Please help me solve this. If you have a recipe that calls for 1 1/4 cups of water to 3 cups of flour, How many cups of water would you need for 2 cups of flour.
----
Use a proportion:::
x/(2 cups of flour) = (5/4)/(3 cups of flour)
----
x = 2[(5/12]
x = 5/6 cups of water
========================
Cheers,
Stan H.
=================